# Classical facets of quantum integrability

### v1

@@
This paper is a review of the works devoted to understanding and reinterpretation of the theory of quantum integrable models solvable by Bethe ansatz in terms of the theory of purely classical soliton equations. Remarkably, studying polynomial solutions of the latter by methods of the classical soliton theory, one is able to develop a method to solve the spectral problem for the former which is alternative to the Bethe ansatz procedure. Our main examples are the generalized inhomogeneous spins chains with twisted boundary conditions from the quantum side and the modified Kadomtsev-Petviashvili hierarchy of nonlinear differential-difference equations from the classical side. In this paper, we restrict ourselves by quantum spin chains with rational $GL(n)$-invariant $R$-matrices (of the XXX type). Also, the connection of quantum spin chains with classical soliton equations implies a close interrelation between the spectral problem for spin chains and integrable many-body systems of classical mechanics such as Calogero-Moser and Ruijsenaars-Scheider models, which is known as the quantum-classical duality. Revisiting this topic, we suggest a simpler and more instructive proof of this kind of duality.
@@ 1 Introduction
An intriguing phenomenon in the world of integrable models, which already has been observed on many examples, is that the models belonging to very different classes (for example, such as mechanical systems, spin chains, nonlinear partial differential equations) have nontrivial hidden interrelations. One of such kind of interrelations is existence of various dualities connecting differently looking models. Another one is a rather mysterious appearance of classical integrable equations as exact relations built in the structure of quantum integrable systems even at $\hbar\neq 0$. Moreover, such relations sometimes allow one to develop alternative methods to diagonalize commuting Hamiltonians of the quantum systems, based on purely classical theory.
The latter program was first realized in [2] (see also [3, 4]), where functional relations for higher conserved quantities of generalized quantum spin chains [5, 6, 7] were interpreted as a classical discrete dynamical system which was identified with the Hirota bilinear difference equation [8] known in soliton theory since 1981 (it is often referred to as a fully discrete KP equation). It was demonstrated that this interpretation allows one to solve the spectral problem for spin chains using methods of classical soliton theory. Later, in [9, 10] this approach was extended to spin chains with superalgebra symmetries (graded magnets).
@@
where $s_{\lambda}({\bf t})$ are Schur polynomials [25] and the sum goes over all Young diagrams including the empty one (actually, for $GL(n)$-invariant models the sum is restricted to diagrams with not more than $n$ non-empty rows). In [11] the generating function (1.3) was called the master $T$-operator (see also [26], where it was introduced in an implicit form). The main its property, proved in [11], is that it satisfies the bilinear equations for the tau-function of the mKP hierarchy, with $x$ playing the role of the “zeroth time” variable. The generating bilinear equation has the form [27, 28]
@@
A further analysis shows that the objects arising in the process of solving the quantum model using the algebraic Bethe ansatz technique and specific for the quantum theory have their counterparts in the classical theory of soliton equations. For example, the classical facet of the nested Bethe ansatz procedure is a chain of Bäcklund transformations (we call it the “undressing chain”) which gradually “undress” a solution of the mKP hierarchy to the trivial one. Each step of this chain corresponds to a level of the nested Bethe ansatz. The eigenvalues of the Baxter’s $Q$-operators are nothing else than tau-functions arising at different steps of the undressing chain. Then the Bethe equations for their roots acquire a nice interpretation as equations of motion for the Ruijsenaars-Schneider system of particles in discrete time (each step of the discrete time corresponds to a level of the Bethe ansatz). Further, the non-commutative generating function for transfer matrices in fundamental representations becomes, on the classical side, the so-called wave operator ${\bf W}(x)$ (known also as the dressing operator). In general, it is an infinite series in inverse powers of the shift operator $e^{\eta\partial_{x}}$ but a characteristic feature of solutions relevant to the $GL(n)$-invariant quantum spin chains is that this series truncates at the $n$-th term, and this operator admits a factorization in a finite product of first order difference operators. This factorization plays a key role in the algebraic Bethe ansatz solution since the coefficients of the first order operators in each factor are constructed from eigenvalues of the $Q$-operators $Q_{k}(x)$, $k=1,\ldots,n-1$, so eigenvalues of the transfer matrices can be expressed through the $Q_{k}$’s whose roots are subject to the system of Bethe equations.
@@
The dynamics of zeros of polynomial tau-functions is a well known subject in the theory of integrable nonlinear partial differential equations. In the works by Krichever and others (see [29][36]) it was found that this dynamics is described by equations of motion of integrable many-body systems of particles of the Calogero-Moser and Ruijsenaars-Schneider type. In particular, the dynamics of zeros of the tau-function of the mKP hierarchy of the form (1.5) in the time $t_{k}$ coincides with the dynamics of the Ruijsenaars-Schneider system of particles [37] (which is also known as a relativistic deformation of the Calogero-Moser system [38, 39], the parameter $\eta$ playing the role of the inverse velocity of light) with respect to the $k$-th Hamiltonian flow (for all $k$ this result was obtained in [40]). The Ruijsenaars-Schneider system admits the Lax representation with the Lax matrix
@@
where $M_{a}$ are eigenvalues of the weight operators ${{\bf M}_{a}=\sum_{k=1}^{N}\pi_{(1)}({\bf e}^{(k)}_{aa})}$ on the common eigenstate with the transfer matrix and $p_{i}$ are the twist parameters (elements of the diagonal twist matrix ${\bf g}$).
@@
The structure of the paper is as follows. In Section 2 we recall the standard definitions and facts from the theory of integrable $GL(n)$-invariant spin chains. In Section 3 we introduce the higher transfer matrices ${\sf T}_{\lambda}(x)$ indexed by the Young diagrams $\lambda$ and recall the functional relations for them, which have the form of the spectral parameter dependent determinant identities of the Jacobi-Trudi type. The master $T$-operator is introduced in Section 4.1, and in Section 4.2 the integral bilinear relation for it is given. This relation allows one to identify eigenvalues of the master $T$-operator with tau-functions of the mKP hierarchy. Section 5 is devoted to the basic facts related to the mKP hierarchy. In particular, we recall the construction of the wave operator ${\bf W}(x,{\bf t})$, the wave function $\psi(x,{\bf t};z)$ and its adjoint $\psi^{*}(x,{\bf t};z)$. In Section 6 we study (quasi)polynomial solutions to the mKP hierarchy. They are constructed using the approach suggested by Krichever, which is based on the conditions (6.2) imposed to the wave function as a function of the spectral parameter $z$. Also, in this section the undressing and dressing Bäcklund transformations are considered, and a chain of the Bäcklund transformations is introduced. The index that labels steps of the chain can be regarded as a discrete time variable. The dynamics of zeros of the tau-functions in this discrete time is the Ruijsenaars-Schneider system in discrete time. Remarkably, its equations of motion are the nested Bethe equations. In Section 7 we present a detailed identification of the objects from the algebraic Bethe ansatz with the ones from the classical theory of the mKP hierarchy. In Section 8 the connection with the classical Ruijsenaars-Schneider many-body system is established, and the quantum-classical duality is discussed. Finally, in Section 9 we make a few concluding remarks and mention some open problems.
@@ 2.1 $GL(n)$-invariant $R$-matrices
Let $V_{1},V_{2},V_{3}$ be three copies of the space $V$. By ${\sf R}_{12}(x)$ we denote the $R$-matrix that acts nontrivially in $V_{1}\otimes V_{2}$ and trivially in $V_{3}$, etc. The $R$-matrix is required to satisfy the Yang-Baxter equation
@@
where the both sides are linear operators in the space $V_{1}\otimes V_{2}\otimes V_{3}$. The $GL(n)$-invariance means that the ${\sf R}(x)$ commutes with ${\bf g}\otimes{\bf g}$ for any ${\bf g}\in GL(n)$, i.e.,
@@
where ${\sf I}$ is the unity matrix in $V\otimes V$, $e_{ab}$ are elementary $n\!\times\!n$ matrices with 1 at the place $ab$ and 0 otherwise and $\eta$ is a parameter [^1] This parameter can be absorbed in $x$ but we prefer to keep it alive in order be prepared for the limit to the Gaudin model which is the limit $\eta\to 0$.. Note that
@@ 2.2 Transfer matrices
where trace is taken in the auxiliary space $V_{0}$. The transfer matrix is a linear operator in the space ${\cal V}$. The Yang-Baxter equation for the $R$-matrix guarantees that the transfer matrices commute at different values of the spectral parameter: $[{\sf T}(x),\,{\sf T}(x^{\prime})]=0.$
@@
where both sides are operators in ${\cal V}=V_{1}\otimes V_{2}\otimes\ldots\otimes V_{N}$ and ${\bf g}_{j}$ acts as ${\bf g}$ in $V_{j}$ and trivially in the other tensor factors.
@@ 3 Transfer matrices as generalized characters
Two simple special cases are important for what follows. The empty diagram $\lambda=\emptyset$ corresponds to the trivial representation: $\pi_{\emptyset}({\bf e}_{ab})=0$. The column of $n$ boxes corresponds to one-dimensional representation $\pi_{(1^{n})}({\bf e}_{ab})=\delta_{ab}$.
@@ 3.1 Higher transfer matrices
It is known that the characters are given by Schur polynomials $s_{\lambda}$ of eigenvalues $g_{i}$ of the matrix ${\bf g}$:
@@
The Schur polynomials corresponding to the diagrams which are rows or columns of the form, respectively, $\lambda=(s)$ or $\lambda=1^{a}$ play especially important role. It is customary to use the special notation for them [25]:
@@
where the sum in the left-hand side is taken over all Young diagrams including the empty one. Second, there are identities which express Schur polynomials for general $\lambda$’s through those corresponding to the diagrams which are either rows or columns, i.e., through $h_{\lambda}({\bf t})$ or $e_{\lambda}({\bf t})$:
@@
The transfer matrices in symmetric or antisymmetric representations, i.e., corresponding to the diagrams which are rows or columns of the form, respectively, $\lambda=(s)$ or $\lambda=1^{a}$ play especially important role. In what follows we will use the special simplified notation for them:
@@
In fact the relations (3.19) and (3.20) are equivalent: (3.20) follows from (3.19) and vice versa. Note that since ${\sf T}_{\lambda}(x)$ are polynomials, the possible poles at $x=x_{i}\pm k\eta$ coming from the pre-factors in the right-hand sides, must cancel by zeros of the determinants. Like the characters, the transfer matrices ${\sf T}_{\lambda}(x)$ vanish identically if $\lambda_{1}^{\prime}>n$ (in particular, ${\sf T}^{a}(x)$ vanishes if $a>n$).
@@ 4.2 Bilinear functional relations for the master $T$-operator
Note that there is a freedom to multiply the tau-function by $A^{x}$, where $A$ is any non-zero constant: the transformation $c_{\lambda}(x)\to A^{x}c_{\lambda}(x)$ does not spoil the relations (4.9), (4.10). We call the tau-functions which differ by a factor of the form $A^{x}$ equivalent.
As it follows from the bilinear formalism [27, 28], the generating bilinear relation for the master $T$-operator can be written in the form
@@ 6.3 Undressing transformations
$\begin{array}[]{l}\displaystyle{\tau^{(m)}(x,{\bf t})=\mathop{\hbox{res}}\limits_{z_{j}=p_{j}\atop j=m+1,\ldots,n}\left[\prod_{\alpha=m+1}^{n}(z_{\alpha}^{-x/\eta+m-n}e^{\xi({\bf t},z_{\alpha})})\Delta_{m}(z_{m+1},\ldots,z_{n})\right.}\\ \\ \phantom{aaaaaaaaaaaaaaa}\displaystyle{\left.\times\,\tau\Bigl{(}x+(n-m)\eta,{\bf t}+\!\!\sum_{\alpha=m+1}^{n}[z_{\alpha}^{-1}]\Bigr{)}\right],}\end{array}$ (6.27)
where ${\Delta_{n}(z_{1},\ldots,z_{m})=\prod\limits_{i>j}^{m}(z_{i}-z_{j})}$ is the Vandermonde determinant. In particular, on the previous to the last level we have
@@ 9 Concluding remarks
Finally, let us recall that the quantum-classical duality receives a formal “explanation” if one treats it as a limiting case of the so-called Matsuo-Cherednik correspondence, which was established in [47, 48] (see also [49, 50], where it was generalized and connected with the quantum-classical duality). The Matsuo-Cherednik correspondence, sometimes called the quantum-quantum duality, connects solutions to the (quantum) Knizhnik-Zamolodchikov equations (which can be regarded as a non-stationary extension of the spectral problem for the spin chain Hamiltonians ${\bf H}_{j}$) with stationary wave functions of the quantized Ruijsenaars-Schneider model. The parameter that controls non-stationarity in the former plays the role of the Planck’s constant in the latter. Tending it to zero makes the former problems stationary but still quantum, while the latter models become classical, and this is the way how the quantum-quantum (Matsuo-Cherednik) duality turns to the quantum-classical duality. As we have seen, the quantum-classical duality can be lifted to the level of integrable hierarchies of nonlinear equations like KP or mKP. It is then natural to ask whether something similar is true for the quantum-quantum duality, i.e., could it be obtained from a more general connection with an integrable hierarchy of KP or mKP type (probably, quantized).

### v2

@@
This paper is a review of the works devoted to understanding and reinterpretation of the theory of quantum integrable models solvable by Bethe ansatz in terms of the theory of purely classical soliton equations. Remarkably, studying polynomial solutions of the latter by methods of classical soliton theory, one is able to develop a method of solving the spectral problem for the former which provides an alternative to the Bethe ansatz procedure. Our main examples are the generalized inhomogeneous spins chains with twisted boundary conditions on the quantum side and the modified Kadomtsev-Petviashvili hierarchy of nonlinear differential-difference equations on the classical side. In this paper, we restrict ourselves to quantum spin chains with rational $GL(n)$-invariant $R$-matrices (of the XXX type). Also, the connection of quantum spin chains with classical soliton equations implies a close interrelation between the spectral problem for spin chains and integrable many-body systems of classical mechanics such as Calogero-Moser and Ruijsenaars-Scheider models, which is known as the quantum-classical duality. Revisiting this topic, we suggest a simpler and more instructive proof of this kind of duality.
@@ 1 Introduction
An intriguing phenomenon in the world of integrable models, which already has been observed in many examples, is that the models belonging to very different classes (for example, such as mechanical systems, spin chains, nonlinear partial differential equations) have nontrivial hidden interrelations. One such interrelation is the existence of various dualities connecting differently looking models. Another one is a rather mysterious appearance of classical integrable equations as exact relations built into the structure of quantum integrable systems even at $\hbar\neq 0$. Moreover, such relations sometimes allow one to develop alternative methods to diagonalize commuting Hamiltonians of the quantum systems, based on purely classical theory.
The latter program was first realized in [2] (see also [3, 4]), where functional relations for higher conserved quantities of generalized quantum spin chains [5, 6, 7] were interpreted as a classical discrete dynamical system which was identified with the Hirota bilinear difference equation [8] known in soliton theory since 1981. (It is often referred to as a fully discrete KP equation.) It was demonstrated that this interpretation allows one to solve the spectral problem for spin chains using methods of classical soliton theory. Later, in [9, 10] (see also [11] for a further development) this approach was extended to generalized spin chains with superalgebra symmetries (super-spin chains or “graded magnets” introduced by in [12]). This extension was based on the functional relations for higher transfer matrices of the super-spin chains established in [13, 14].
@@
where $s_{\lambda}({\bf t})$ are Schur polynomials [29] and the sum goes over all Young diagrams including the empty one (actually, for $GL(n)$-invariant models the sum is restricted to diagrams with not more than $n$ non-empty rows). In [15] the generating function (1.3) was called the master $T$-operator (see also [30], where it was introduced in an implicit form). Its main property, proved in [15], is that it satisfies the bilinear equations for the tau-function of the mKP hierarchy, with $x$ playing the role of the “zeroth time” variable. The generating bilinear equation has the form [31, 32]
@@
A further analysis shows that the objects arising in the process of solving the quantum model using the algebraic Bethe ansatz technique and specific for the quantum theory have their counterparts in the classical theory of soliton equations. For example, the classical facet of the nested Bethe ansatz procedure is a chain of Bäcklund transformations (which we call the “undressing chain”) that gradually “undress” a solution of the mKP hierarchy to the trivial one. Each step of this chain corresponds to a level of the nested Bethe ansatz. The eigenvalues of the Baxter’s $Q$-operators are nothing else than tau-functions arising at different steps of the undressing chain. Then the Bethe equations for their roots acquire a nice interpretation of equations of motion for the Ruijsenaars-Schneider system of particles in discrete time (each step of the discrete time corresponds to a level of the Bethe ansatz). Further, the non-commutative generating function for transfer matrices in fundamental representations becomes, on the classical side, the so-called wave operator ${\bf W}(x)$ (known also as the dressing operator). In general, it is an infinite series in inverse powers of the shift operator $e^{\eta\partial_{x}}$ but a characteristic feature of solutions relevant to the $GL(n)$-invariant quantum spin chains is that this series truncates at the $n$-th term, and this operator admits a factorization as a finite product of first order difference operators. This factorization plays a key role in the algebraic Bethe ansatz solution since the coefficients of the first order operators in each factor are constructed from eigenvalues of the $Q$-operators $Q_{k}(x)$, $k=1,\ldots,n-1$, so eigenvalues of the transfer matrices can be expressed through the $Q_{k}$’s whose roots are subject to the system of Bethe equations.
@@
The dynamics of the zeros of polynomial tau-functions is a well known subject in the theory of integrable nonlinear partial differential equations. In the works of Krichever and others (see [33][40]) it was found that these dynamics are described by equations of motion of integrable many-body systems of particles of the Calogero-Moser and Ruijsenaars-Schneider type. In particular, the dynamics of zeros of the tau-function of the mKP hierarchy of the form (1.5) in the times $t_{k}$ coincide with the dynamics of the Ruijsenaars-Schneider system of particles [41] (which is also known as a relativistic deformation of the Calogero-Moser system [42, 43], the parameter $\eta$ playing the role of the inverse velocity of light) with respect to the $k$-th Hamiltonian flow (for all $k$ this result was obtained in [44]). The Ruijsenaars-Schneider system admits the Lax representation with the Lax matrix
@@
where $M_{a}$ are the eigenvalues of the weight operators ${{\bf M}_{a}=\sum_{k=1}^{N}\pi_{(1)}({\bf e}^{(k)}_{aa})}$ on the common eigenstate with the transfer matrix and $p_{i}$ are the twist parameters (elements of the diagonal twist matrix ${\bf g}$).
@@
The structure of the paper is as follows. In Section 2 we recall the standard definitions and facts from the theory of integrable $GL(n)$-invariant spin chains. In Section 3 we introduce the higher transfer matrices ${\sf T}_{\lambda}(x)$ indexed by the Young diagrams $\lambda$ and recall the functional relations for them, which have the form of the spectral parameter dependent determinant identities of the Jacobi-Trudi type. The master $T$-operator is introduced in Section 4.1, and in Section 4.2 the integral bilinear relation for it is given. This relation allows one to identify eigenvalues of the master $T$-operator with tau-functions of the mKP hierarchy. Section 5 is devoted to the basic facts related to the mKP hierarchy. In particular, we recall the construction of the wave operator ${\bf W}(x,{\bf t})$, the wave function $\psi(x,{\bf t};z)$ and its adjoint $\psi^{*}(x,{\bf t};z)$. In Section 6 we study (quasi)polynomial solutions to the mKP hierarchy. They are constructed using the approach suggested by Krichever, which is based on the conditions (6.2) imposed on the wave function as a function of the spectral parameter $z$. Also, in this section the undressing and dressing Bäcklund transformations are considered, and a chain of the Bäcklund transformations is introduced. The index that labels steps of the chain can be regarded as a discrete time variable. The dynamics of zeros of the tau-functions in this discrete time is the Ruijsenaars-Schneider system in discrete time. Remarkably, its equations of motion are the nested Bethe equations. In Section 7 we present a detailed identification of the objects from the algebraic Bethe ansatz with the ones from the classical theory of the mKP hierarchy. In Section 8 the connection with the classical Ruijsenaars-Schneider many-body system is established, and the quantum-classical duality is discussed. Finally, in Section 9 we make a few concluding remarks and mention some open problems.
@@ 2.1 $GL(n)$-invariant $R$-matrices
Let $V_{1},V_{2},V_{3}$ be three copies of the space $V$. By ${\sf R}_{12}(x)$ we denote the $R$-matrix that acts nontrivially on $V_{1}\otimes V_{2}$ and trivially in $V_{3}$, etc. The $R$-matrix is required to satisfy the Yang-Baxter equation
@@
where the both sides are linear operators on the space $V_{1}\otimes V_{2}\otimes V_{3}$. The $GL(n)$-invariance means that the ${\sf R}(x)$ commutes with ${\bf g}\otimes{\bf g}$ for any ${\bf g}\in GL(n)$, i.e.,
@@
where ${\sf I}$ is the unity matrix in $V\otimes V$, $e_{ab}$ are elementary $n\!\times\!n$ matrices with 1 at the place $ab$ and 0 otherwise and $\eta$ is a parameter [^1] This parameter can be absorbed in $x$ but we prefer to keep it alive in order to prepare for the limit to the Gaudin model which is $\eta\to 0$.. Note that
@@ 2.2 Transfer matrices
where the trace is taken in the auxiliary space $V_{0}$. The transfer matrix is a linear operator in the space ${\cal V}$. The Yang-Baxter equation for the $R$-matrix guarantees that the transfer matrices commute at different values of the spectral parameter: $[{\sf T}(x),\,{\sf T}(x^{\prime})]=0.$
@@
where both sides are operators on ${\cal V}=V_{1}\otimes V_{2}\otimes\ldots\otimes V_{N}$ and ${\bf g}_{j}$ acts as ${\bf g}$ in $V_{j}$ and trivially in the other tensor factors.
@@ 3 Transfer matrices as generalized characters
Two simple special cases are important for what follows. The empty diagram $\lambda=\emptyset$ corresponds to the trivial representation: $\pi_{\emptyset}({\bf e}_{ab})=0$. The column of $n$ boxes corresponds to the one-dimensional representation $\pi_{(1^{n})}({\bf e}_{ab})=\delta_{ab}$.
@@ 3.1 Higher transfer matrices
It is known that the characters are given by Schur polynomials $s_{\lambda}$ of the eigenvalues $g_{i}$ of the matrix ${\bf g}$:
@@
The Schur polynomials corresponding to the diagrams that are rows or columns of the form, respectively, $\lambda=(s)$ or $\lambda=1^{a}$ play a specially important role. It is customary to use the special notation for them [29]:
@@
where the sum on the left-hand side is taken over all Young diagrams including the empty one. Second, there are identities which express Schur polynomials for general $\lambda$’s through those corresponding to the diagrams which are either rows or columns, i.e., through $h_{\lambda}({\bf t})$ or $e_{\lambda}({\bf t})$:
@@
The transfer matrices in symmetric or antisymmetric representations, i.e., corresponding to the diagrams which are rows or columns of the form, respectively, $\lambda=(s)$ or $\lambda=1^{a}$ play a specially important role. In what follows we will use the special simplified notation for them:
@@
In fact the relations (3.19) and (3.20) are equivalent: (3.20) follows from (3.19) and vice versa. Note that since ${\sf T}_{\lambda}(x)$ are polynomials, the possible poles at $x=x_{i}\pm k\eta$ coming from the pre-factors in the right-hand sides, must be cancelled by zeros of the determinants. Like the characters, the transfer matrices ${\sf T}_{\lambda}(x)$ vanish identically if $\lambda_{1}^{\prime}>n$ (in particular, ${\sf T}^{a}(x)$ vanishes if $a>n$).
@@ 4.2 Bilinear functional relations for the master $T$-operator
Note that there is a freedom to multiply the tau-function by $A^{x}$, where $A$ is any non-zero constant: the transformation $c_{\lambda}(x)\to A^{x}c_{\lambda}(x)$ does not spoil the relations (4.9), (4.10). We call the tau-functions that differ by a factor of the form $A^{x}$ equivalent.
It follows from the bilinear formalism [31, 32] that the generating bilinear relation for the master $T$-operator can be written in the form
@@ 6.3 Undressing transformations
$\begin{array}[]{l}\displaystyle{\tau^{(m)}(x,{\bf t})=\mathop{\hbox{res}}\limits_{z_{j}=p_{j}\atop j=m+1,\ldots,n}\left[\prod_{\alpha=m+1}^{n}(z_{\alpha}^{-x/\eta+m-n}e^{\xi({\bf t},z_{\alpha})})\Delta(z_{m+1},\ldots,z_{n})\right.}\\ \\ \phantom{aaaaaaaaaaaaaaa}\displaystyle{\left.\times\,\tau\Bigl{(}x+(n-m)\eta,{\bf t}+\!\!\sum_{\alpha=m+1}^{n}[z_{\alpha}^{-1}]\Bigr{)}\right],}\end{array}$ (6.27)
where ${\Delta(x_{1},\ldots,x_{m})=\prod\limits_{i>j}^{m}(x_{i}-x_{j})}$ is the Vandermonde determinant. In particular, on the previous to the last level we have
@@ 9 Concluding remarks
Finally, let us recall that the quantum-classical duality receives a formal “explanation” if one treats it as a limiting case of the so-called Matsuo-Cherednik correspondence, which was established in [51, 52] (see also [53, 54], where it was generalized and connected with the quantum-classical duality). The Matsuo-Cherednik correspondence, sometimes called the quantum-quantum duality, connects solutions to the (quantum) Knizhnik-Zamolodchikov equations (which can be regarded as a non-stationary extension of the spectral problem for the spin chain Hamiltonians ${\bf H}_{j}$) with stationary wave functions of the quantized Ruijsenaars-Schneider model. The parameter that controls non-stationarity in the former plays the role of the Planck’s constant in the latter. Tending it to zero makes the former problems stationary but still quantum, while the latter models become classical, and this is the way how the quantum-quantum (Matsuo-Cherednik) duality turns to the quantum-classical duality. As we have seen, the quantum-classical duality can be lifted to the level of integrable hierarchies of nonlinear equations like KP or mKP. It is then natural to ask whether something similar is true for the quantum-quantum duality, i.e., could it be obtained from a more general connection with an integrable hierarchy of KP or mKP type or its quantized version.